The perimeter of a rectangle is 30m and the length is 10m more than the width.find the dimensions of the rectangle
Answer by macston(5194)

P=2(L+W)=30 m; L=W+10
30 m=2(W+10+W)
30 m=4W+20 subtract 20 from each side
30m-20m=4W+20m-20m
10m=4W Divide each side by 4
10 m/4=4W/4
2.5 m=W width=2.5 m
L=W+10=12.5 Length = 12.5 m
CHECK
P=2(L+W)
30 m=2(12.5m+2.5m)
30 m=2(15m)
30 m=30 m